OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = -625 and a2 = 125? (1 points) -0.2 0.04 -0.04 0.2

OpenStudy (campbell_st):

well start by finding the common ratio.... r it can be found by using \[r = \frac{a_{2}}{a_{1}}\]

rishavraj (rishavraj):

find the common ratio first of all...nd then find the seventh term

OpenStudy (anonymous):

Ok So Its Going To Be r= 125/-625 @campbell_st

OpenStudy (campbell_st):

well you can simplify that fraction... then use the formula you have been given 2 or 3 times before for the sum of a geometric series

OpenStudy (anonymous):

I Got 0.2

OpenStudy (campbell_st):

ok... oops so a term in a series is \[a_{n} = a_{1} \times r^{n - 1}\] you know the 1st term, common ratio r and n = 7 substitute and then calculate

OpenStudy (campbell_st):

and you value of r should be r = -0.2

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so its going to be \[-625\times \left( -0.2 \right)^n-1\] But wat is n

OpenStudy (anonymous):

@campbell_st